OpenStudy (anonymous):

An artificial satellite circles the Earth in a circular orbit at a location where the acceleration due to gravity is 4.26 m/s2. Determine the orbital period of the satellite.

OpenStudy (anonymous):

Note: \(a=\frac{v^2}{r}\) and \(v=\frac{2 \pi}{T}\) do you need further help?